# Data curation, Methodology, Writing - original draft

### v3

@@ 1 Introduction
Medical image segmentation plays a crucial role in numerous clinical applications, such as computer-aided diagnoses [1][2], image-guided interventions [3][4], [5][6][7], and surgical robotics [8][9][10][11]. UNet [12] and its variants [13][14][15][16] have been widely used in this field, achieving tremendous success in different medical imaging modalities. However, accurate and efficient segmentation of 3D medical images still remains a non-trivial task [17].
@@
Current mainstream segmentation methods can be classified into two categories: 2D-based and 3D-based methods [18]. 2D-based methods split 3D images into 2D slices and segment them individually, while 3D-based methods divide 3D images into smaller patches and then segment these patches individually. Despite impressive performance achieved by state-of-the-art methods [19], they still exhibit some limitations. Most 2D-based methods focus on architecture design to enhance intra-slice representations for better performance, such as incorporating attention modules [20][21] or adopting transformers [22][23]. However, these methods overlook inter-slice cues, which are also crucial for accurate segmentation. In contrast, 3D-based methods can capture intra- and inter-slice information for segmentation but demand substantial GPU memory and computational resources. Additionally, they tend to perform poorly in images with anisotropic voxel spacing since they are primarily designed for 3D images with nearly isotropic voxel spacing [24][25].
@@
In order to combine the advantages of 2D-based and 3D-based methods, some studies have been done to explore 2.5D-based segmentation models [18]. The main idea of these methods is to fuse inter-slice (neighborhood slices) information into 2D-based models when segmenting specific slices (target slices). The most direct way to achieve inter-slice fusion is by concatenating slices as multi-channel inputs. However, it is inefficient, making it challenging for models to extract useful features for the target slice [18]. Therefore, some studies focus on exploring “smart” ways of inter-slice fusion. Most of them formulate 2D slices as time sequences and adopt recurrent neural network (RNN) [26], transformers [27][28] or attention mechanisms [29] to fuse inter-slice information.
@@
While current 2.5D-based methods have achieved impressive segmentation results, they struggle with distinguishing individual slices during inter-slice fusion, and consequently fail to learn reliable inter-slice representations essential for accurate segmentation [18]. The root of this problem lies in the use of a single encoder to process all input slices, resulting in the same feature distributions across the feature space, as shown in Fig. 1 (a). For example, the features of the $i$-th slice remain identical whether it is considered as the target slice or the neighborhood slice. This indistinguishability becomes problematic in scenarios where consecutive slices, such as the $i$-th and the $(i+1)$-th, are target slices, respectively. Models fail to differentiate the $i$-th slice’s features as belonging to the target or the neighborhood slice, thereby hampering the extraction of valuable inter-slice information for segmentation.
@@
To address the above issue, a novel 2.5D-based segmentation model, MOSformer, MOmentum encoder-based inter-Slice fusion transformer is proposed to effectively leverage inter-slice information for 3D medical image segmentation. MOSformer follows the design of the U-shaped architecture [12]. In order to enhance feature distinguishability of each slice, dual encoders are utilized in our model, with one for target slices and the other for neighborhood slices. Parameters of the target slice encoder are updated by back-propagation, and parameters of the neighborhood slice encoder are updated using a momentum update. Therefore, features can remain distinguishable and consistent, promoting inter-slice fusion, as shown in Fig. 1 (b). Furthermore, we propose an efficient inter-slice fusion Transformer (IF-Trans) that captures inter-slice cues from multi-scale feature maps at each scale, built upon Swin Transformer [30].
@@ 2 Related Works
Several 2.5D-based approaches have been proposed for efficient medical image segmentation by leveraging inter-slice information. Early methods concatenated multiple consecutive 2D slices into a multi-channel input and adopted 2D-based models to segment specific regions of the middle slice [31][32]. However, concatenating 2D slices as multi-channel inputs hinders the model’s capacity to disentangle and learn slice-specific features [18], thereby constraining the performance of 2.5D-based models. To solve the above problem, some works treated continuous 2D slices as temporal sequences and utilized recurrent neural networks (RNNs) [26][33] to learn inter-slice information. For example, Chen et al. [26] introduced a 2.5D segmentation framework that combines $k$-UNet and bi-directional convolutional LSTM (BDC-LSTM) to integrate inter-slice information. Although RNNs can help improve the performance of 2.5D-based models to some extent, training costs of these methods are considerably high [34]. Instead of RNNs, recent studies utilized attention mechanisms or transformers to fuse inter-slice information at the feature level effectively. Zhang et al. [29] proposed an attention fusion module to refine segmentation results by fusing the information of adjacent slices. Li et al. [35] employed a 2.5D coarse-to-fine architecture that leveraged inter-slice prediction discrepancies as spatial attention cues to refine the initial segmentation. Guo et al. [36] adopted 2D UNet as the backbone and fused inter-slice information via a transformer at the bottom layer of the encoder. Yan et al. [27] proposed AFTer-UNet with an axial fusion mechanism based on transformer to fuse intra- and inter-slice contextual information. Hung et al. [28][37] and Kumar et al. [38] introduced novel cross-slice attention mechanisms based on transformer to learn cross-slice information at multiple scales. However, the aforementioned methods fail to capture useful inter-slice information for the target slice which needs to be segmented, since they use a single encoder to extract slice features, making it difficult for models to distinguish target slices from neighborhood slices [18].
@@ 2.2 Transformers in Medical Image Segmentation
Recently, with the tremendous success of vision transformer (ViT) [39] in various computer vision tasks [40][41], many works have explored using transformers in medical image segmentation. Compared with CNNs, transformers can capture long-range dependencies by sequence modeling and multihead self-attention (MHSA) [39], achieving better segmentation performance. Chen et al. [23] proposed TransUNet, combining UNet [12] and transformer, where the transformer encodes feature maps from the CNN encoder to extract global contexts for the decoder to generate segmentation results. Cao et al. [22] are the first to employ a pure transformer architecture for medical image segmentation. Convolutional layers in UNet are all replaced by blocks in Swin transformer [30]. However, this architecture does not achieve better performance [42]. Huang et al. introduced MISSformer [43], which incorporates an encoder-decoder architecture built on enhanced transformer blocks. These blocks are connected through the ReMixed transformer context bridge, enhancing the model’s ability to capture discriminative details. You et al. [42] presented CASTformer with a class-aware transformer module to better capture discriminative regions of target objects. Moreover, they utilized adversarial learning to boost segmentation accuracies. However, the 2D-based methods mentioned above face limitations in leveraging inter-slice information, which hinders their potential for further performance improvements. Some attempts have been made to build 3D-based transformer segmentation models. UNETR [44] pioneered the use of a transformer-based encoder to learn global contexts from volumetric data. CoTr [45] introduced a deformable self-attention mechanism to reduce computational complexity. However, simplifying self-attention may cause contextual information loss [27]. nnFormer [46] is an interleaved architecture, where convolution layers encode precise spatial information and transformer layers fully explore global dependencies. Similar to Swin transformer [30], a computationally efficient way to calculate self-attention is proposed in nnFormer.
@@
In this work, we introduce the inter-slice fusion transformer (IF-Trans), which extends Swin Transformer’s (shifted) window multi-head self-attention [30], (S)W-MSA, to fuse inter-slice information. Unlike prior methods that restrict transformer attention to each 2D slice, IF-Trans lifts (S)W-MSA into the inter-slice domain. Each window attends not only to patches within its own slice but also to the corresponding windows in neighborhood slices. This design preserves Swin Transformer’s computational efficiency while jointly modeling intra-slice and inter-slice context, improving 3D medical image segmentation.
@@ 3.1 Overall Architecture
The detailed architecture of MOSformer is shown in Fig. 2. Like most previous works for medical image segmentation, we utilize a hybrid encoder-decoder architecture, combining the advantages of CNNs and transformers [46]. $\bm{x}_{i}\in\mathbb{R}^{C\times H\times W}$ is the input of the encoder and represents the target slice for segmentation, where $i$ indicates the $i$-th slice of a 3D volume $\bm{X}\in\mathbb{R}^{C\times H\times W\times D}$, where $C$, $H$, $W$, and $D$ denote the channels, height, width, and depth of $\bm{X}$. $\bm{x}_{j}\in\mathbb{R}^{C\times H\times W}$ denotes an input to the momentum encoder corresponding to a neighborhood slice of $\bm{x}_{i}$, where $j\in\left[i-s,i+s\right]\setminus\left\{i\right\}$. The hyperparameter $s$ represents the $s$-th neighborhood of $\bm{x}_{i}$. Finally, the model generates the segmentation map $\bm{y}_{i}\in\mathbb{R}^{C_{0}\times H\times W}$ of $\bm{x}_{i}$, where $C_{0}$ is the number of label classes.
We adopt a lightly modified ResNet-50 [47] to implement the encoders, which extract multi-scale features from input slices. Concretely, the final downsampling stage is replaced with a non-downsampling stage to preserve spatial resolution. Dual encoders with a momentum update are adopted in MOSformer to strengthen feature distinguishability and maintain feature consistency. Furthermore, IF-Trans modules are used at multi-scale ($1/2$, $1/4$, $1/8$, and $1/16$) to fuse inter-slice features extracted by dual encoders (details are provided in Section 3.3). Then the fused features are sent to the decoder via skip connections. The final segmentation predictions are derived via a segmentation head ($1\times 1$ convolutional layer).
@@ 3.2 Dual Encoders with A Momentum Update
Conventional 2.5D-based methods employ a single encoder to process both the target slice and its neighborhood slices, then fuse their features at a later stage. However, because all slices share the same feature space, the model struggles to distinguish target-specific cues from neighborhood context [18], as illustrated in Fig. 1 (a). Consequently, inter-slice fusion may be suboptimal, and fine-grained spatial cues of the target slice can be lost. An intuitive solution is to use two separate encoders to process neighborhood slices and target slices, respectively. In practice, however, updating the parameters of these encoders independently during training leads to inconsistent feature distributions, which again hampers effective fusion.
@@
To address this, we draw inspiration from the momentum contrast framework (MoCo) [48] and introduce a momentum encoder for the neighborhood slices. The key idea is to maintain a slowly updating copy of the target encoder, ensuring that neighborhood features remain consistent over training while still being distinguishable from the target features.
@@ 3.3 Inter-slice Fusion Transformer
Different from standard self-attention [49] with quadratic complexity, the proposed IF-Trans only calculates self-attention within the local window. As shown in the left part of Fig. 3, feature maps are partitioned into several non-overlapping windows [^1] For intuitive explanation, feature maps are replaced by input images, and the number of feature pixels is simplified to 16.. Compared with Swin transformer [30], we compute self-attention within inter- and intra-slice local windows (i.e., CSW-MSA, cross-slice window-based multi-head self-attention) instead of intra-slice local windows. Consequently, the purple feature pixel in the target slice attends not only to the yellow pixels within its own slice but also to the green pixels from neighborhood slices, thereby capturing both intra-slice and inter-slice context, as illustrated by the red arrows in Fig. 3.
However, the local CSW-MSA lacks connections across windows, reducing its representational power. Similar to [30], a shifted window partitioning strategy is introduced, allowing each pixel to receive broader views from intra- and inter-slices. In Fig. 3, the first transformer module adopts a regular window partition approach, and the feature map is evenly divided into $2\times 2$ windows of size $2\times 2$ (M=2M=2) [^2] To correspond with Fig. 3, $M$ is set to 2 here.. The second transformer module uses a different partitioning configuration. Windows of the preceding layer are displaced by $\left(\lfloor\frac{M}{2}\rfloor,\lfloor\frac{M}{2}\rfloor\right)$ pixels to generate new windows. By doing so, the orange pixel can conduct self-attention (i.e., SCSW-MSA, a shifted version of CSW-MSA) with more pixels, thereby boosting its representational capacity. In practice, these two configurations are served as two consecutive layers to get an IF-Trans module. Outputs of IF-Trans can be formulated as:
@@
where $\hat{\bm{f}}^{k,l}_{i}$ and $\bm{f}^{k,l}_{i}$ represents output feature maps of the (S)CSW-MSA module $\mathcal{T}^{\rm(S)}$ and the multilayer perceptron (MLP) module $\mathcal{M}$ in the $l$-th layer, respectively. LN indicates layer normalization. The query-key-value (QKV) self-attention [49] in (S)CSW-MSA is computed as follows:

### v4

@@ 1 Introduction
Medical image segmentation plays a crucial role in numerous clinical applications, such as computer-aided diagnoses [zhou2020real][antonelli2022medical], image-guided interventions [li2021unified, li2021real, han2024respiratory, huang2025real, chen2025gvm], and surgical robotics [zhang2024ai, li2023three, zhou2022learning, zhang2025novel]. UNet [ronneberger2015u] and its variants [zhou2019unet++, huang2020unet, ni2022surginet, huang2024spironet] have been widely used in this field, achieving tremendous success in different medical imaging modalities. However, accurate and efficient segmentation of 3D medical images still remains a non-trivial task [tajbakhsh2020embracing].
@@
Current mainstream segmentation methods can be classified into two categories: 2D-based and 3D-based methods [zhang2022bridging]. 2D-based methods split 3D images into 2D slices and segment them individually, while 3D-based methods divide 3D images into smaller patches and then segment these patches individually. Despite impressive performance achieved by state-of-the-art methods [azad2024medical], they still exhibit some limitations. Most 2D-based methods focus on architecture design to enhance intra-slice representations for better performance, such as incorporating attention modules [mou2021cs2][roy2022lwmla] or adopting transformers [cao2022swin][chen2024transunet]. However, these methods overlook inter-slice cues, which are also crucial for accurate segmentation. In contrast, 3D-based methods can capture intra- and inter-slice information for segmentation but demand substantial GPU memory and computational resources. Additionally, they tend to perform poorly in images with anisotropic voxel spacing since they are primarily designed for 3D images with nearly isotropic voxel spacing [cciccek20163d][yu2017automatic].
@@
In order to combine the advantages of 2D-based and 3D-based methods, some studies have been done to explore 2.5D-based segmentation models [zhang2022bridging]. The main idea of these methods is to fuse inter-slice (neighborhood slices) information into 2D-based models when segmenting specific slices (target slices). The most direct way to achieve inter-slice fusion is by concatenating slices as multi-channel inputs. However, it is inefficient, making it challenging for models to extract useful features for the target slice [zhang2022bridging]. Therefore, some studies focus on exploring “smart” ways of inter-slice fusion. Most of them formulate 2D slices as time sequences and adopt recurrent neural network (RNN) [chen2016combine], transformers [yan2022after][hung2022cat] or attention mechanisms [zhang2020sau] to fuse inter-slice information.
@@
While current 2.5D-based methods have achieved impressive segmentation results, they struggle with distinguishing individual slices during inter-slice fusion, and consequently fail to learn reliable inter-slice representations essential for accurate segmentation [zhang2022bridging]. The root of this problem lies in the use of a single encoder to process all input slices, resulting in the same feature distributions across the feature space, as shown in Fig. 1 (a). For example, the features of the $i$-th slice remain identical whether it is considered as the target slice or the neighborhood slice. This indistinguishability becomes problematic in scenarios where consecutive slices, such as the $i$-th and the $(i+1)$-th, are target slices, respectively. Models fail to differentiate the $i$-th slice’s features as belonging to the target or the neighborhood slice, thereby hampering the extraction of valuable inter-slice information for segmentation.
@@
To address the above issue, a novel 2.5D-based segmentation model, MOSformer, MOmentum encoder-based inter-Slice fusion transformer is proposed to effectively leverage inter-slice information for 3D medical image segmentation. MOSformer follows the design of the U-shaped architecture [ronneberger2015u]. In order to enhance feature distinguishability of each slice, dual encoders are utilized in our model, with one for target slices and the other for neighborhood slices. Parameters of the target slice encoder are updated by back-propagation, and parameters of the neighborhood slice encoder are updated using a momentum update. Therefore, features can remain distinguishable and consistent, promoting inter-slice fusion, as shown in Fig. 1 (b). Furthermore, we propose an efficient inter-slice fusion Transformer (IF-Trans) that captures inter-slice cues from multi-scale feature maps at each scale, built upon Swin Transformer [liu2021swin].
@@ 2 Related Works
Several 2.5D-based approaches have been proposed for efficient medical image segmentation by leveraging inter-slice information. Early methods concatenated multiple consecutive 2D slices into a multi-channel input and adopted 2D-based models to segment specific regions of the middle slice [yu2018recurrent][li2021acenet]. However, concatenating 2D slices as multi-channel inputs hinders the model’s capacity to disentangle and learn slice-specific features [zhang2022bridging], thereby constraining the performance of 2.5D-based models. To solve the above problem, some works treated continuous 2D slices as temporal sequences and utilized recurrent neural networks (RNNs) [chen2016combine][yang2018towards] to learn inter-slice information. For example, Chen et al. [chen2016combine] introduced a 2.5D segmentation framework that combines $k$-UNet and bi-directional convolutional LSTM (BDC-LSTM) to integrate inter-slice information. Although RNNs can help improve the performance of 2.5D-based models to some extent, training costs of these methods are considerably high [pang2019deep]. Instead of RNNs, recent studies utilized attention mechanisms or transformers to fuse inter-slice information at the feature level effectively. Zhang et al. [zhang2020sau] proposed an attention fusion module to refine segmentation results by fusing the information of adjacent slices. Li et al. [li2021learning] employed a 2.5D coarse‑to‑fine architecture that leveraged inter‑slice prediction discrepancies as spatial attention cues to refine the initial segmentation. Guo et al. [guo2021transformer] adopted 2D UNet as the backbone and fused inter-slice information via a transformer at the bottom layer of the encoder. Yan et al. [yan2022after] proposed AFTer-UNet with an axial fusion mechanism based on transformer to fuse intra- and inter-slice contextual information. Hung et al. [hung2022cat][hung2024csam] and Kumar et al. [kumar2024flexible] introduced novel cross-slice attention mechanisms based on transformer to learn cross-slice information at multiple scales. However, the aforementioned methods fail to capture useful inter-slice information for the target slice which needs to be segmented, since they use a single encoder to extract slice features, making it difficult for models to distinguish target slices from neighborhood slices [zhang2022bridging].
@@ 2.2 Transformers in Medical Image Segmentation
Recently, with the tremendous success of vision transformer (ViT) [dosovitskiy2020image] in various computer vision tasks [carion2020end][lijin2024dual], many works have explored using transformers in medical image segmentation. Compared with CNNs, transformers can capture long-range dependencies by sequence modeling and multihead self-attention (MHSA) [dosovitskiy2020image], achieving better segmentation performance. Chen et al. [chen2024transunet] proposed a hybrid model, TransUNet, combining UNet [ronneberger2015u] and transformer, where the transformer encodes feature maps from the CNN encoder to extract global contexts for the decoder to generate segmentation results. To fully unleash the transformer’s potential, subsequent research focused on pure transformer architectures. A key challenge is the high computational complexity of self-attention on high-resolution medical images. Swin Transformer [liu2021swin], with its efficient window-based attention, offered a viable solution. Building on this, Cao et al. [cao2022swin] introduced Swin-Unet, the first pure transformer for medical image segmentation, which replaces all convolutions in U-Net with Swin Transformer blocks. However, this architecture does not achieve better performance than hybrid models [you2022class]. Huang et al. introduced MISSformer [huang2023missformer], which incorporates an encoder-decoder architecture built on enhanced transformer blocks. These blocks are connected through the ReMixed transformer context bridge, enhancing the model’s ability to capture discriminative details. You et al. [you2022class] presented CASTformer with a class-aware transformer module to better capture discriminative regions of target objects. Moreover, they utilized adversarial learning to boost segmentation accuracies. However, the 2D-based methods mentioned above face limitations in leveraging inter-slice information, which hinders their potential for further performance improvements. Some attempts have been made to build 3D-based transformer segmentation models. UNETR [hatamizadeh2022unetr] pioneered the use of a transformer-based encoder to learn global contexts from volumetric data. CoTr [xie2021cotr] introduced a deformable self-attention mechanism to reduce computational complexity. However, simplifying self-attention may cause contextual information loss [yan2022after]. nnFormer [zhou2023nnformer] is an interleaved architecture, where convolution layers encode precise spatial information and transformer layers fully explore global dependencies. Similar to Swin Transformer [liu2021swin], a computationally efficient way to calculate self-attention is proposed in nnFormer.
@@
In this work, we introduce the inter-slice fusion transformer (IF-Trans), which extends Swin Transformer’s (shifted) window multi-head self-attention [liu2021swin], (S)W-MSA, to fuse inter-slice information. Unlike prior methods that restrict transformer attention to each 2D slice, IF-Trans lifts (S)W-MSA into the inter-slice domain. Each window attends not only to patches within its own slice but also to the corresponding windows in neighborhood slices. This design preserves Swin Transformer’s computational efficiency while jointly modeling intra-slice and inter-slice context, improving 3D medical image segmentation.
@@ 3.1 Overall Architecture
The detailed architecture of MOSformer is shown in Fig. 2. Like most previous works for medical image segmentation, we utilize a hybrid encoder-decoder architecture, combining the advantages of CNNs and transformers [zhou2023nnformer]. $\bm{x}_{i}\in\mathbb{R}^{C\times H\times W}$ is the input of the encoder and represents the target slice for segmentation, where $i$ indicates the $i$-th slice of a 3D volume $\bm{X}\in\mathbb{R}^{C\times H\times W\times D}$, where $C$, $H$, $W$, and $D$ denote the channels, height, width, and depth of $\bm{X}$. $\bm{x}_{j}\in\mathbb{R}^{C\times H\times W}$ denotes an input to the momentum encoder corresponding to a neighborhood slice of $\bm{x}_{i}$, where $j\in\left[i-s,i+s\right]\setminus\left\{i\right\}$. The hyperparameter $s$ represents the $s$-th neighborhood of $\bm{x}_{i}$. Finally, the model generates the segmentation map $\bm{y}_{i}\in\mathbb{R}^{C_{0}\times H\times W}$ of $\bm{x}_{i}$, where $C_{0}$ is the number of label classes.
We adopt a lightly modified ResNet-50 [he2016deep] to implement the encoders, which extract multi-scale features from input slices. Concretely, the final downsampling stage is replaced with a non-downsampling stage to preserve spatial resolution. Dual encoders with a momentum update are adopted in MOSformer to strengthen feature distinguishability and maintain feature consistency. Furthermore, IF-Trans modules are used at multi-scale ($1/2$, $1/4$, $1/8$, and $1/16$) to fuse inter-slice features extracted by dual encoders (details are provided in Section 3.3). Then the fused features are sent to the decoder via skip connections. The final segmentation predictions are derived via a segmentation head ($1\times 1$ convolutional layer).
@@ 3.2 Dual Encoders with A Momentum Update
Conventional 2.5D-based methods employ a single encoder to process both the target slice and its neighborhood slices, then fuse their features at a later stage. However, because all slices share the same feature space, the model struggles to distinguish target-specific cues from neighborhood context [zhang2022bridging], as illustrated in Fig. 1 (a). Consequently, inter-slice fusion may be suboptimal, and fine-grained spatial cues of the target slice can be lost. An intuitive solution is to use two separate encoders to process neighborhood slices and target slices, respectively. In practice, however, updating the parameters of these encoders independently during training leads to inconsistent feature distributions, which again hampers effective fusion.
@@
To address this, we draw inspiration from the momentum contrast framework (MoCo) [he2020momentum] and introduce a momentum encoder for the neighborhood slices. The key idea is to maintain a slowly updating copy of the target encoder, ensuring that neighborhood features remain consistent over training while still being distinguishable from the target features.
@@ 3.3 Inter-slice Fusion Transformer
Different from standard self-attention [Vaswani2017attn] with quadratic complexity, the proposed IF-Trans only calculates self-attention within the local window. As shown in the left part of Fig. 3, feature maps are partitioned into several non-overlapping windows [^1] For intuitive explanation, feature maps are replaced by input images, and the number of feature pixels is simplified to 16.. Compared with Swin Transformer [liu2021swin], we compute self-attention within inter- and intra-slice local windows (i.e., CSW-MSA, cross-slice window-based multi-head self-attention) instead of intra-slice local windows. Consequently, the purple feature pixel in the target slice attends not only to the yellow pixels within its own slice but also to the green pixels from neighborhood slices, thereby capturing both intra-slice and inter-slice context, as illustrated by the red arrows in Fig. 3.
However, the local CSW-MSA lacks connections across windows, reducing its representational power. Similar to [liu2021swin], a shifted window partitioning strategy is introduced, allowing each pixel to receive broader views from intra- and inter-slices. In Fig. 3, the first transformer module adopts a regular window partition approach, and the feature map is evenly divided into $2\times 2$ windows of size $2\times 2$ (M=2M=2) [^2] To correspond with Fig. 3, $M$ is set to 2 here.. The second transformer module uses a different partitioning configuration. Windows of the preceding layer are displaced by $\left(\lfloor\frac{M}{2}\rfloor,\lfloor\frac{M}{2}\rfloor\right)$ pixels to generate new windows. By doing so, the orange pixel can conduct self-attention (i.e., SCSW-MSA, a shifted version of CSW-MSA) with more pixels, thereby boosting its representational capacity. In practice, these two configurations are served as two consecutive layers to get an IF-Trans module. Outputs of IF-Trans can be formulated as:
@@
where $\hat{\bm{f}}^{k,l}_{i}$ and $\bm{f}^{k,l}_{i}$ represents output feature maps of the (S)CSW-MSA module $\mathcal{T}^{\rm(S)}$ and the multilayer perceptron (MLP) module $\mathcal{M}$ in the $l$-th layer, respectively. LN indicates layer normalization. The query-key-value (QKV) self-attention [Vaswani2017attn] in (S)CSW-MSA is computed as follows: